(a)  Given the formula:-


Moles  = Concentration(Moldm-3) X Volume(dm3)

            

Use the formula to find the number of moles in


2.5 X 10 -6 dm3 of a 8.0 X 10 -6 moldm-3 solution of potassium   permanganate.



(b)           Evaluate showing your working out without calculator use:-     


           2.6x107 x 3x10-4      

                1.3 x 103

   


1
Expert's answer
2021-03-09T05:22:31-0500

(a)

"8.0 \times 10^{-6} mol\cdot dm^{-3}\times2.5\times 10^{-6}dm^3"

"=2\times10^{-11} mol"

(b)


"2.6 \times 10^{7}\times 3 \times 10^{-4} \times 1.3 \times 10^{3}"

"=1.014\times10^7"